Hartford,Connecticut and Topeka, Kansas are roughly 1260 miles apart. If it took 3 hours and 20 minutes for a plane to fly into the wind from Hartford to Topeka, and 2 hours for a plane to make the return trip with the wind, what was the air speed of the plane and what was the speed of the wind??
Answer by TimothyLamb(4379)

s = d / t
---
outbound:
s - w = 1260/(3+1/3)
s - w = 1260/(10/3)
s - w = 378
s = 378 + w
---
return:
s + w = 1260/2
s + w = 630
---
s + w = 630
378 + w + w = 630
378 + 2w = 630
2w = 252
---
w = 126 mph
s = 504 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php